OpenStudy (sarahhanish):

What is the slope of the line represented by the equation 2x+3y= –12 GIVE ME THE ANSWER THIS IS THE LAST ONE YESSS ? –32 –23 23 32

OpenStudy (sarahhanish):

CAN U GIVE IT TO ME IM DESPRETE

OpenStudy (sarahhanish):

yessss people looking at thisss

OpenStudy (steve816):

Again, you need to get the equation into y = mx+b form.

OpenStudy (sarahhanish):

bu-but i dont even know what that even means lol

OpenStudy (3mar):

\[\Large 2x+3y=-12\] \[\Large 2x+3y-2x=-12-2x\] \[\Large 3y=-12-2x\] \[\Large 3y=-2x-12\] \[\Large \frac{ 3y }{ 3 } =\frac{ -2x-12 }{ 3 }\] \[\Large y =-\frac{ 2x }{ 3 }-\frac{ 12 }{ 3 }\] \[\Huge\color{DeepPink}{y =-\frac{ 2 }{ 3 }x-4}\] \[\Huge\color{DeepPink}{y =~~~m~x+b}\]

OpenStudy (sarahhanish):

WOAH I JUST WANTED THE ANSWER LOL

OpenStudy (sarahhanish):

SO WHATS THE ANSWER

OpenStudy (steve816):

-2/3

OpenStudy (steve816):

Remember that m is the slope!

OpenStudy (sarahhanish):

OH

OpenStudy (3mar):

It is not helpful for you to be given the answers without understanding how they are figured out!

OpenStudy (3mar):

Compare the two pink equations, you will pick what the slope is! @sarahhanish

OpenStudy (sarahhanish):

THE FIRST?

OpenStudy (3mar):

\[\Huge\color{DeepPink}{y =-\frac{ 2 }{ 3 }x-4}\] \[\Huge\color{DeepPink}{y =~~~m~x+b}\] so m=?? and b=??

OpenStudy (3mar):

@sarahhanish Did you get it?